OpenStudy (anonymous):

Find the horizontal, vertical, and oblique asymptotes, if any, for the given rational function G(x)=(x^4-1)/(3x^2-3x)

OpenStudy (anonymous):

\[ g(x)=\frac{x^4-1}{3x^2-3x}\\ g(x)=\frac{(x^2+1)(x^2-1)}{3x(x-1)}\\ g(x)=\frac{(x^2+1)(x+1)\cancel{(x-1)}}{3x\cancel{(x-1)}} \]

OpenStudy (anonymous):

we can cancel only if \(x\ne1\) 1) vertical asymptote: the value of "x" for which the function becomes undefined. this is a rational function. When the denominator becomes "0", we cannot solve it, what value of "x" makes it 0? 2) horizontal asymptote: rearrange the equation interms of "y" and see if you will get a rational function or "x" becomes undefined for some value of "y"

OpenStudy (anonymous):

So the vertical is x=0 and the horizontal is y = (1/3)(x^2 + x + 1) ?

OpenStudy (anonymous):

3) oblique asymptote: perform the polynomial division: \[g(x)=\frac{(x^2+1)(x+1)}{3x}\]

OpenStudy (anonymous):

how did you get the horizontal asymptote? horizontal asymptote must be a straight line with slope = 0

OpenStudy (anonymous):

What would the horizontal look like?

OpenStudy (anonymous):

(1/3)(x^2 + x + 1) + 0 , which is a paraloba

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

horizontal means Flat. no curve

OpenStudy (anonymous):

so, it should simply be: y = a number

OpenStudy (anonymous):

you still did not tell me how you found \[{1\over 3}(x^2+x+1)\]

OpenStudy (anonymous):

when x --> ∞ 1/3x --> 0 so we left with (1/3)(x^2 + x + 1) + 0 , which is a paraloba by simplifying

OpenStudy (anonymous):

so is the horizontal 1/3?

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

horizontal is y=0

OpenStudy (anonymous):

that parabola is an oblique asymptote. you have to extract two lines from it. and there is no horizontal asymptote

OpenStudy (anonymous):

(1/3)(x^2 + x + 1) + 0 so this is the oblique?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

not this y = (1/3)(x^2 + x + 1) ?